# Approximation in Hölder spaces

### v2

@@ 3.3. Approximation of Banach-valued mappings from $c_{0}$
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $\varepsilon>0.$ Then, there exists $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $r>0$ so that $\|g-f\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon,$ and for every $x\in X,$ there exists a finite subset $S=S_{x}$ of $\mathcal{A}$ for which $g(z)=g(P_{S}(z)),$ whenever $z\in B(x,r).$ Moreover, if $f$ has bounded support (resp. Lipschitz), then the above $g$ has bounded supported (resp. Lipschitz) too.
@@
We have shown that $\|f-g\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon.$ Now, if $x\in X=c_{0}(\mathcal{A}),$ then (by definition) there exists a finite subset $S=S_{x}$ such that if $\alpha\in\mathcal{A}\setminus S,$ then $|x_{\alpha}|\leq r/2.$ Then, for $z\in B(x,r/2)$ and for each $\alpha\in\mathcal{A}\setminus S$ it follows that $|z_{\alpha}|\leq r,$ implying that $\varphi(z_{\alpha})=0.$ It follows for $z\in B(x,r/2)$ that $\phi(z)=\phi(P_{S}(z))$ and hence that $g(z)=g(P_{S}(z)).$ For the second part, suppose there exists $R>0$ so that $f(x)=0$ for all $\|x\|\geq R.$ If $\|z\|\geq R+r,$ then $|z_{\alpha_{0}}|\geq R+r$ for some $\alpha_{0}\in\mathcal{A}.$ Consequently
@@
and thus $g(z)=f(\phi(z))=0.$ Finally, if $f$ is Lipschitz, then clearly $g=f\circ\phi$ is Lipschitz. ∎
@@
In the proof of part (i), by Lemma 3.13, given a fixed $r>0,$ it is enough to approximate functions $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ with the property that for every $x\in X,$ there exists finite $S_{x}\subset\mathcal{A}$ such that $g(z)=g(P_{S_{x}}(z)),$ for all $z\in B(x,2r).$ Given $\varepsilon,$ let $\delta>0$ be so that
@@
Since $z\mapsto h_{S_{x}}(z)$ is of class $C^{\infty}(B(x,r),Y)$, the above yields $h\in C^{\infty}(X,Y).$ All these properties were stated and proved in [HajJoh09, Lemma 6]. We next show that $h\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X,Y).$ Given any two points $x,z\in X,$ let $S_{x},S_{z}\subset\mathcal{A}$ be the finite subsets associated with $x$ and $z,$ and set $S=S_{x}\cup S_{z}.$ Since $h(x)=h_{S}(x)$ and $h(z)=h_{S}(z)$ we have
@@
We conclude that $\|g-h\|_{\dot{C}^{0,\omega}(X,Y)}\leq 2\varepsilon$ and complete the proof of Theorem 3.12(i). For part (ii), given $f\in\dot{\operatorname{VC}}^{0,\omega}(X,Y),$ by Theorem 1.4, we approximate $f$, in the $\dot{C}^{0,\omega}(X,Y)$-seminorm by a $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ with bounded support. By Lemma 3.13, we can assume that for every $x\in X$ there exists a finite $S=S_{x}$ such that $g=g\circ P_{S}$ on $B(x,r),$ and for a uniform fixed $r>0.$ Let us next check that the approximation $h$ from part (i) has bounded support. Let $R>0$ be so that $g$ is zero on $X\setminus B(0,R).$ Assume that $\|x\|\geq R+\eta.$ There exists a finite set $S=S_{x}\subset\mathcal{A}$ so that $h(x)=h_{S}(x)$. Now, for $t=(t_{\alpha})_{\alpha\in S}\in[-\eta,\eta]^{|S|},$ there holds that
@@
The proof of Theorem 3.12(ii) is now complete. Finally, in the particular case $Y=\mathbb{R},$ let us establish the desired Lipschitz approximations. Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ be the function to be approximated. By Theorem 3.3 we can assume that $f$ is Lipschitz, and by Lemma 3.13, there is a Lipschitz approximation $g$ of $f$, with $g$ locally depending only on a finite number of coordinates around each ball of fixed radius $r>0.$ Repeating the construction of $h\in C^{\infty}(X,)$, we immediately see from the estimate (47), that $h$ is Lipschitz, and in fact $\operatorname{Lip}(h)\leq\operatorname{Lip}(g).$ Because any Lipschitz mapping in $\dot{C}^{0,\omega}(X)$ belongs to $\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X),$ by the condition (3), the identity in $(i)$ now follows at once. As for approximation of an $f\in\dot{\operatorname{VC}}^{0,\omega}(X)$, we apply Theorem 3.3 to reduce matters to $f\in\operatorname{Lip}_{\operatorname{bs}}(X)$. Then the function $g$ of Lemma 3.13 can be taken Lipschitz and with bounded support, and so can the function $h$ we constructed in (i) and (ii) of the present theorem. Thus $h\in C^{\infty}_{\operatorname{bs}}(X)\cap\operatorname{Lip}(X)$ approximates $f$ in the $\dot{C}^{0,\omega}(X)$ seminorm. Because any Lipschitz mapping with bounded support belongs to $\dot{\operatorname{VC}}^{0,\omega}(X),$ the identity in (ii) now follows at once. ∎
@@
Acknowledgements
C. M. was supported by the grant no. 334466 of the Research Council of Norway, “Fourier Methods and Multiplicative Analysis”. T.O. was supported by the Finnish Academy of Sciences and Letters.

### v3

@@ 3.3. Approximation of Banach-valued mappings from $c_{0}$
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $\varepsilon>0.$ Then, there exists $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $r>0$ so that $\|g-f\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon,$ and for every $x\in X,$ there exists a finite subset $S=S_{x}$ of $\mathcal{A}$ for which $g(z)=g(P_{S}(z)),$ whenever $z\in B(x,r).$
Moreover, if $f$ has bounded support (resp. Lipschitz), then the above $g$ has bounded supported (resp. Lipschitz) too.
@@
We have shown that $\|f-g\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon.$
Now, if $x\in X=c_{0}(\mathcal{A}),$ then (by definition) there exists a finite subset $S=S_{x}$ such that if $\alpha\in\mathcal{A}\setminus S,$ then $|x_{\alpha}|\leq r/2.$ Then, for $z\in B(x,r/2)$ and for each $\alpha\in\mathcal{A}\setminus S$ it follows that $|z_{\alpha}|\leq r,$ implying that $\varphi(z_{\alpha})=0.$ It follows for $z\in B(x,r/2)$ that $\phi(z)=\phi(P_{S}(z))$ and hence that $g(z)=g(P_{S}(z)).$
For the second part, suppose there exists $R>0$ so that $f(x)=0$ for all $\|x\|\geq R.$ If $\|z\|\geq R+r,$ then $|z_{\alpha_{0}}|\geq R+r$ for some $\alpha_{0}\in\mathcal{A}.$ Consequently
@@
and thus $g(z)=f(\phi(z))=0.$ Finally, if $f$ is Lipschitz, then clearly $g=f\circ\phi$ is Lipschitz.
@@
In the proof of part (i), by Lemma 3.13, given a fixed $r>0,$ it is enough to approximate functions $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ with the property that for every $x\in X,$ there exists finite $S_{x}\subset\mathcal{A}$ such that $g(z)=g(P_{S_{x}}(z)),$ for all $z\in B(x,2r).$
Given $\varepsilon,$ let $\delta>0$ be so that
@@
Since $z\mapsto h_{S_{x}}(z)$ is of class $C^{\infty}(B(x,r),Y)$, the above yields $h\in C^{\infty}(X,Y).$ All these properties were stated and proved in [HajJoh09, Lemma 6].
We next show that $h\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X,Y).$ Given any two points $x,z\in X,$ let $S_{x},S_{z}\subset\mathcal{A}$ be the finite subsets associated with $x$ and $z,$ and set $S=S_{x}\cup S_{z}.$ Since $h(x)=h_{S}(x)$ and $h(z)=h_{S}(z)$ we have
@@
We conclude that $\|g-h\|_{\dot{C}^{0,\omega}(X,Y)}\leq 2\varepsilon$ and complete the proof of Theorem 3.12(i).
For part (ii), given $f\in\dot{\operatorname{VC}}^{0,\omega}(X,Y),$ by Theorem 1.2, we approximate $f$, in the $\dot{C}^{0,\omega}(X,Y)$-seminorm by a $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ with bounded support. By Lemma 3.13, we can assume that for every $x\in X$ there exists a finite $S=S_{x}$ such that $g=g\circ P_{S}$ on $B(x,r),$ and for a uniform fixed $r>0.$ Let us next check that the approximation $h$ from part (i) has bounded support. Let $R>0$ be so that $g$ is zero on $X\setminus B(0,R).$ Assume that $\|x\|\geq R+\eta.$ There exists a finite set $S=S_{x}\subset\mathcal{A}$ so that $h(x)=h_{S}(x)$. Now, for $t=(t_{\alpha})_{\alpha\in S}\in[-\eta,\eta]^{|S|},$ there holds that
@@
The proof of Theorem 3.12(ii) is now complete.
Finally, in the particular case $Y=\mathbb{R},$ let us establish the desired Lipschitz approximations. Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ be the function to be approximated. By Theorem 3.3 we can assume that $f$ is Lipschitz, and by Lemma 3.13, there is a Lipschitz approximation $g$ of $f$, with $g$ locally depending only on a finite number of coordinates around each ball of fixed radius $r>0.$ Repeating the construction of $h\in C^{\infty}(X,)$, we immediately see from the estimate (43), that $h$ is Lipschitz, and in fact $\operatorname{Lip}(h)\leq\operatorname{Lip}(g).$ Because any Lipschitz mapping in $\dot{C}^{0,\omega}(X)$ belongs to $\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X),$ by the condition (3), the identity in $(i)$ now follows at once.
As for approximation of an $f\in\dot{\operatorname{VC}}^{0,\omega}(X)$, we apply Theorem 3.3 to reduce matters to $f\in\operatorname{Lip}_{\operatorname{bs}}(X)$. Then the function $g$ of Lemma 3.13 can be taken Lipschitz and with bounded support, and so can the function $h$ we constructed in (i) and (ii) of the present theorem. Thus $h\in C^{\infty}_{\operatorname{bs}}(X)\cap\operatorname{Lip}(X)$ approximates $f$ in the $\dot{C}^{0,\omega}(X)$ seminorm. Because any Lipschitz mapping with bounded support belongs to $\dot{\operatorname{VC}}^{0,\omega}(X),$ the identity in (ii) now follows at once.
@@
Appendix A Comparison of $\dot{C}^{0,\omega}$ and uniform convergence
We begin with a remark concerning the two types of convergences we are dealing with. We will refer to the norm $|||\cdot|||_{C^{0,\omega}(X,Y)}$ defined in (1).
Remark A.1.
Let $X$ and $Y$ be normed spaces and $\omega$ any modulus of continuity. Suppose that a sequence of functions $(f_{n})_{n}\subset\dot{C}^{0,\omega}(X,Y)$ converges to $f,$ with respect to the norm $|||\cdot|||_{C^{0,\omega}(X,Y)}.$ Then also $f_{n}\to f$ uniformly on bounded subsets of $X.$
Indeed, let $B$ a bounded subset of $X.$ Given $\varepsilon>0,$ let $N$ be so that $|||f_{n}-f|||_{C^{0,\omega}(X,Y)}\leq\varepsilon/(1+\omega(\operatorname{diam}(B)))$ for all $n\geq N.$ Then, for those $n\geq N,$ and $x\in B,$ we can write
$|f(x)-f_{n}(x)|$ $\leq|f(0)-f_{n}(0)|+|(f-f_{n})(x)-(f-f_{n})(0)|$
$\leq\left(1+\omega(\|x\|)\right)|||f_{n}-f|||_{C^{0,\omega}(X,Y)}\leq\left(1+\omega(\operatorname{diam}(B))\right)|||f_{n}-f|||_{C^{0,\omega}(X,Y)}\leq\varepsilon.$
However, the convergence in $\dot{C}^{0,\omega}$ does not imply global uniform convergence, even if we assume that the pertinent sequence of functions is globally uniformly bounded, as the next example shows.
Example A.2.
Let $\omega$ be of the form $\omega(t)=t^{\alpha},$ with $0<\alpha\leq 1$ Define $f_{n}:\mathbb{R}\to\mathbb{R}$ as the continuous function given by $f_{n}=0$ on $(-\infty,0]\cup[2n,+\infty),$ $f_{n}(n)=1$, and $f_{n}$ affine on both $[0,n]$ and $[n,2n].$ It is immediate that $\|f_{n}\|_{\infty}\leq 1$ for all $n,$ and that $(f_{n})_{n}$ does not converge to 00 uniformly on $\mathbb{R}.$ However, one has $f_{n}\to 0$ in the $\dot{C}^{0,\omega}(\mathbb{R})$ sense. This easily follows after observing that
$\sup_{n\leq s<t\leq 2n}\frac{|f_{n}(t)-f_{n}(s)|}{\omega(|t-s|)}=\sup_{0\leq s<t\leq n}\frac{|f_{n}(t)-f_{n}(s)|}{\omega(|t-s|)}=\sup_{0\leq s<t\leq n}\frac{1}{n}\frac{t-s}{(t-s)^{\alpha}}=\frac{1}{n^{\alpha}}.$
Finally, we next show that the converse to Remark A.1 is false, by providing a sequence $(f_{n})_{n}\subset\dot{C}^{0,\omega}$ and $f\in\dot{C}^{0,\omega}$ so that $f_{n}$ converges uniformly on $\mathbb{R},$ and $f_{n}$ does not converge to $f$ in the $\|\cdot\|_{\dot{C}^{0,\omega}}$ norm on bouded subsets.
Example A.3.
Define $f=0$ on $\mathbb{R}$ and the continuous functions $f_{n}:\mathbb{R}\to[0,1]$ given by
$f_{n}(t)=\begin{cases}1/\sqrt{n},\qquad&t=0,\\ \text{affine},\qquad&t\in[0,1/n],\\ 0,\qquad&t\geq 1/n,\\ f_{n}(-t),\qquad&t\leq 0.\end{cases}$
Then $\sup_{t\in\mathbb{R}}|f_{n}(t)|=1/\sqrt{n},$ yielding $\|f_{n}-f\|_{\infty}\to 0.$ Also, each $f_{n}$ is $\sqrt{n}$-Lipschitz:
$\frac{|f_{n}(0)-f_{n}(1/n)|}{|0-1/n|}=\frac{1/\sqrt{n}}{1/n}=\sqrt{n}.$
Because $f_{n}$ is bounded, then $f_{n}\in\dot{C}^{0,\omega}(\mathbb{R})$ for all moduli $\omega$ that satisfy $\lim_{t\to 0}t/\omega(t)<\infty.$
However, if $\omega(t)=t^{\alpha}$ with $\alpha\in(1/2,1],$ then $\lim_{n}\|f_{n}-f\|_{\dot{C}^{0,\omega}([0,1])}=\infty,$ since
$\lim_{n}\|f_{n}-f\|_{\dot{C}^{0,\omega}([0,1])}\geq\lim_{n}\frac{|f_{n}(0)-f_{n}(1/n)|}{\omega(|0-1/n|)}=\lim_{n}n^{\alpha-\frac{1}{2}}=\infty.$